# DiffusionNFT: Online Diffusion Reinforcement with Forward Process

### v1

@@ 1 Introduction
Despite promising efforts made, we argue that GRPO-style diffusion reinforcement still faces fundamental limitations: (1) Forward inconsistency. Focusing solely on the reverse sampling process breaks adherence to the forward diffusion process, risking the model degenerating into cascaded Gaussians. (2) Solver restriction. The training loss is entangled with first-order SDE samplers, precluding the full utilization of ODE or high-order solvers that are default to flow models and advantageous for generation efficiency. (3) Complicated CFG integration. Diffusion models heavily rely on Classifier-Free Guidance (CFG) (Ho & Salimans, 2022), which requires training both conditional and unconditional models. Current RL practices directly incorporate CFG in post-training, leading to a complicated and inefficient two-model optimization scheme.
There is a single forward (noising) but various reverse (denoising) processes for a diffusion policy. Can diffusion reinforcement be performed on the forward process instead of the reverse?
We propose a new online RL paradigm: Diffusion Negative-aware FineTuning (DiffusionNFT). DiffusionNFT is not built on the Policy Gradient framework. Instead, it performs policy optimization directly on the forward diffusion process through the flow matching objective. Intuitively, DiffusionNFT splits generated samples into positive and negative subsets and defines a contrastive improvement direction between the two policies learned on them. Moreover, unlike conventional guidance techniques that modify the sampling process, DiffusionNFT adopts an implicit parameterization technique that allows integrating reinforcement guidance directly into the optimized policy.
@@
The forward-process RL formulation provides several practical benefits (Figure 2). First, DiffusionNFT allows data collection with arbitrary black-box solvers, rather than just first-order SDE samplers. Second, it eliminates the need to store entire sampling trajectories, requiring only clean images for policy optimization. Third, it is fully compatible with standard diffusion training, requiring minimal modifications to existing codebases. Finally, it is a native off-policy algorithm, naturally allowing decoupled training and sampling policies without importance sampling.
@@
Our main contribution is DiffusionNFT, a direct RL alternative to conventional Policy Gradient methods, introducing the Negative-aware FineTuning (NFT) paradigm (Chen et al., 2025c) into the diffusion domain. Grounded in a supervised learning foundation, we believe this paradigm offers a valid path toward a general, unified, and native off-policy RL recipe across various modalities.
@@ 2.1 Diffusion and Flow Models
Rectified flow (Liu et al., 2022) can be considered as a simplified special case of the above-discussed diffusion models, where $\alpha_{t}=1-t,\sigma_{t}=t$, which simplifies the velocity target to ${\bm{v}}=\bm{\epsilon}-{\bm{x}}_{0}$.
@@ 3.1 Problem Setup
Consider a pretrained diffusion policy $\pi^{\text{old}}$ and prompt datasets $\{{\bm{c}}\}$. At each iteration, we sample $K$ images ${\bm{x}}_{0}^{1:K}$ for prompt $\bm{c}$}, and then evaluate each image with a scalar reward function $r\in[0,1]$, representing the its optimality probability $r({\bm{x}}_{0},{\bm{c}}):=p(\mathbf{o}=1|{\bm{x}}_{0},{\bm{c}})$ (Levine, 2018).
Collected data can be randomly split into two imaginary subsets. An image ${\bm{x}}_{0}$ will have a probability $r$ of falling into the positive dataset $\mathcal{D}^{+}$ and otherwise the negative dataset $\mathcal{D}^{-}$. Given infinite samples, the underlying distributions of these two subsets are respectively
@@
It is easy to prove that $\pi^{+}\succ\pi^{\text{old}}\succ\pi^{-}$ constantly holds, thus a straightforward improvement of $\pi^{\text{old}}$ can be $\pi^{*}=\pi^{+}$. To achieve this, previous work (Lee et al., 2023) performs diffusion training solely on $\mathcal{D}^{+}$, known as Rejection FineTuning (RFT). Despite the simplicity, RFT cannot effectively leverage negative data in $\mathcal{D}^{-}$ (Chen et al., 2025c).
@@
We posit that negative feedback is crucial to policy improvement. Rather than treating $\pi^{+}$ as an optimization point, we leverage both negative and positive data to derive an improvement direction $\Delta\in\mathbb{R}^{n}$. The training target is defined as
@@ 3.2 Negative-aware Diffusion Reinforcement with Forward Process
Having defined a valid optimization target $\bm{v}$$^{*}$ with Eq. (3) and 3.1, we now introduce a training objective that directly optimizes ${\bm{v}}_{\theta}$ towards $\bm{v}$$^{*}$:
@@
1. Forward Consistency. In contrast to policy gradient methods (e.g., FlowGRPO), which formulated RL on the reverse diffusion process, DiffusionNFT defines a typical diffusion loss on the forward process. This preserves what we term forward consistency—the adherence of the diffusion model’s underlying probability density to the Fokker-Planck equation (Øksendal, 2003; Song et al., 2020b), ensuring that the learned model corresponds to a valid forward process rather than degenerating into cascaded Gaussians.
@@ 3.3 Practical Implementation
Adaptive Loss Weighting. Typical diffusion loss includes a time-dependent weighting $w(t)$ (Eq. (1)). Instead of manual tuning, we adopt an adaptive weighting scheme. The velocity predictor ${\bm{v}}_{\theta}$ can be equivalently transformed into ${\bm{x}}_{0}$ predictor, denoted as ${\bm{x}}_{\theta}$ (e.g., ${\bm{x}}_{\theta}={\bm{x}}_{t}-t{\bm{v}}_{\theta}$ under rectified flow schedule). We replace the weighting with a form of self-normalized ${\bm{x}}_{0}$ regression:
@@
where sg is the stop-gradient operator. We find it offers stable training (Figure 9).
CFG-Free Optimization. Classifier-Free Guidance (CFG) (Ho & Salimans, 2022) is a default technique to enhance generation quality at inference time, yet it complicates post-training and reduces efficiency. Conceptually, we interpret CFG as an offline form of reinforcement guidance (Eq. (3.1)), where conditional and unconditional models correspond to positive and negative signals. With this understanding, we discard CFG in our algorithm design. The policy is initialized solely by the conditional model. Despite this seemingly poor initialization, we observe that performance surges and quickly surpasses the CFG baseline (Figure 1). This suggests that the functionality of CFG can be effectively learned or substituted through RL post-training, echoing recent studies that achieve strong performance without CFG through post-training (Chen et al., 2025b; a; Zheng et al., 2025).
@@ 5 Related Work
Likelihood-based methods: (1) Diffusion-DPO (Wallace et al., 2024; Yang et al., 2024; Liang et al., 2024; Yuan et al., 2024; Li et al., 2025a) adapts DPO to diffusion for paired human preference data but requires additional likelihood and loss approximations compared to AR. (2) Policy gradient methods, starting from PPO style (Black et al., 2023; Fan et al., 2023), decompose trajectory likelihoods step by step without considering forward consistency. Recent GRPO extensions (Liu et al., 2025; Xue et al., 2025) prove effective and scalable for diffusion RL, but they couple the training loss with SDE samplers and face efficiency bottlenecks. MixGRPO (Li et al., 2025b) improves efficiency by mixing SDE and ODE, while issues of coupling and forward inconsistency remain.
@@ 6 Conclusion
We introduce Diffusion Negative-aware FineTuning (DiffusionNFT), a new paradigm for online reinforcement learning of diffusion models that directly operates on the forward process. By formulating policy improvement as a contrast between positive and negative generations, DiffusionNFT integrates reinforcement signals seamlessly into the standard diffusion objective, eliminating the reliance on likelihood estimation and SDE-based reverse process. Empirically, DiffusionNFT demonstrates strong and efficient reward optimization, achieving up to $25\times$ higher efficiency than FlowGRPO while producing a single omni-model that outperforms CFG baselines across diverse in-domain and out-of-domain rewards. We believe this work represents a step toward unifying supervised and reinforcement learning in diffusion, and highlights the forward process as a promising foundation for scalable, efficient, and theoretically principled diffusion RL.
@@ B.1 Flow SDE
To leverage the diffusion SDE formulation in Song et al. (2020b), we need to match its forward SDE $\mathrm{d}{\bm{x}}_{t}=f(t){\bm{x}}_{t}\mathrm{d}t+g(t)\mathrm{d}{\bm{w}}_{t}$ with the forward transition kernel ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}\bm{\epsilon}$. As noted in the first two versions of VDM (Kingma et al., 2021), $f(t),g(t)$ are related to $\alpha_{t},\sigma_{t}$ by $f(t)=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}$, $g^{2}(t)=\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma_{t}^{2}$. Setting $\alpha_{t}=1-t,\sigma_{t}=t$, we have
@@
where $\lambda_{t}\in[0,1]$. Equivalently, it amounts to introducing Langevin dynamics on top of the diffusion ODE, with $\lambda_{t}=0$ corresponding to ODE, and $\lambda_{t}=1$ corresponding to the maximum variance SDE in Song et al. (2020b). The score function ${\bm{s}}_{\theta}({\bm{x}}_{t},t)\approx\nabla_{{\bm{x}}_{t}}\log\pi_{t}({\bm{x}}_{t})$, noise predictor $\bm{\epsilon}_{\theta}({\bm{x}}_{t},t)$, data predictor ${\bm{x}}_{\theta}({\bm{x}}_{t},t)$ and velocity predictor ${\bm{v}}_{\theta}({\bm{x}}_{t},t)$ are interconvertible (Zheng et al., 2023b):
@@
where $\rho_{t}=\eta_{t}s\sqrt{1-\frac{s^{2}(1-t)^{2}}{t^{2}(1-s)^{2}}}$, and $\eta_{t}\in[0,1]$ interpolates between ODE and maximum variance SDE. Compared to the Euler discretization, the DDIM-style discretization avoids singularities at boundaries and is expected to reduce sampling errors. However, as the training loss of FlowGRPO couples with the sampler, and RL requires not only accurate reward signals but also large exploration space, we did not observe notable advantages by replacing the SDE sampler with DDIM.
@@ B.3 Intuition behind the FlowGRPO Objective
Therefore, FlowGRPO essentially aligns the velocity field with the advantage-weighted noise, while the choice of timesteps and sampler only influences the weighting $\frac{m(s,t)}{n(s,t)}$ across sampling steps. From this perspective, the training loss of FlowGRPO should be disentangled with sampler parameters like the variance $g_{t}$, which current practices fail to address.

### v2

@@ 1 Introduction
Despite promising efforts made, we argue that GRPO-style diffusion reinforcement still faces fundamental limitations: (1) Forward inconsistency. Focusing solely on the reverse sampling process breaks adherence to the forward diffusion process, risking the model degenerating into cascaded Gaussians. (2) Solver restriction. The data collection process relies on first-order SDE samplers, precluding the full utilization of ODE or high-order solvers that are default to flow models and advantageous for generation efficiency. (3) Complicated CFG integration. Diffusion models heavily rely on Classifier-Free Guidance (CFG) (Ho & Salimans, 2022), which requires training both conditional and unconditional models. Current RL practices typically incorporate CFG in post-training, leading to a complicated and inefficient two-model optimization scheme.
We aim to disentangle data collection, remove solver restriction, and maintain consistency with standard supervised pretraining in diffusion RL. As a diffusion policy admits a single forward (noising) process but multiple reverse (denoising) processes (e.g., different samplers), a natural question is:
Can diffusion reinforcement be performed on the forward process instead of the reverse?
This paper proposes a novel online RL paradigm named Diffusion Negative-aware FineTuning (DiffusionNFT). Instead of building upon the conventional policy gradient framework, DiffusionNFT directly performs policy optimization on the forward diffusion process through the flow matching objective. Intuitively, it defines a contrastive improvement direction between two implicit policies learned on “positive” and “negative” generated samples split by reward signals, and optimizes toward the positive policy without modifying the sampling process.
@@
The forward-process RL formulation provides several practical benefits (Figure 2). First, DiffusionNFT allows data collection with arbitrary black-box solvers, rather than relying on first-order SDE samplers. Second, it eliminates the need to store entire sampling trajectories, requiring only clean images for policy optimization. Third, it is fully compatible with standard diffusion training, requiring minimal modifications to existing codebases. Finally, it is a native off-policy algorithm, naturally allowing decoupled training and sampling policies without importance sampling.
@@
DiffusionNFT is a direct RL alternative to conventional Policy Gradient methods, introducing the Negative-aware FineTuning (NFT) paradigm (Chen et al., 2025c) into the diffusion domain. Grounded in a supervised learning foundation, we believe this paradigm offers a valid path toward a general, unified, and native off-policy RL recipe across various modalities.
@@ 2.1 Diffusion and Flow Models
Rectified flow (Liu et al., 2022) can be considered as a special case of the above-discussed diffusion models, where $\alpha_{t}=1-t,\sigma_{t}=t$, which simplifies the velocity target to ${\bm{v}}=\bm{\epsilon}-{\bm{x}}_{0}$.
@@ 3.1 Problem Setup
Consider a pretrained diffusion policy $\pi^{\text{old}}$ and prompt datasets $\{{\bm{c}}\}$. At each iteration, we sample $K$ images ${\bm{x}}_{0}^{1:K}$ for prompt $\bm{c}$}, and then evaluate each image with a scalar reward function $r\in[0,1]$, representing its optimality probability $r({\bm{x}}_{0},{\bm{c}}):=p(\mathbf{o}=1|{\bm{x}}_{0},{\bm{c}})$ (Levine, 2018).
This optimality serves as a bridge from continuous-valued rewards to a binary partition. Collected data can be randomly split into two imaginary subsets. An image ${\bm{x}}_{0}$ will have a probability $r$ of falling into the positive dataset $\mathcal{D}^{+}$ and otherwise the negative dataset $\mathcal{D}^{-}$. Given infinite samples, the underlying distributions of these two subsets are respectively
@@
It is easy to prove that $\pi^{+}\succ\pi^{\text{old}}\succ\pi^{-}$ constantly holds, thus a straightforward improvement of $\pi^{\text{old}}$ can be $\pi^{*}=\pi^{+}$. To achieve this, previous work (Lee et al., 2023) performs diffusion training solely on $\mathcal{D}^{+}$, known as Rejection FineTuning (RFT).
Despite the simplicity, RFT cannot effectively leverage negative data in $\mathcal{D}^{-}$ (Chen et al., 2025c).
@@
We posit that negative feedback is crucial to policy improvement, especially for diffusion [^1] We find that finetuning only on the positive data leads to collapse (Section 4.4). Rather than treating $\pi^{+}$ as an optimization point, we leverage both negative and positive data to derive an improvement direction $\Delta\in\mathbb{R}^{n}$. The training target is defined as
@@ 3.2 Negative-aware Diffusion Reinforcement with Forward Process
Having defined a valid optimization target $\bm{v}$$^{*}$ with Eq. (3) and (3.1), we now introduce a training objective that directly optimizes ${\bm{v}}_{\theta}$ towards $\bm{v}$$^{*}$:
@@
1. Forward Consistency. In contrast to policy gradient methods (e.g., FlowGRPO), which formulated RL on the reverse diffusion process, DiffusionNFT defines a typical diffusion loss on the forward process. This preserves what we term forward consistency—the adherence of the diffusion model’s underlying probability density to the Fokker-Planck equation (Øksendal, 2003; Song et al., 2020b), ensuring that the learned model corresponds to a valid forward process (i.e., ${\bm{x}}_{t}$ are correctly coupled with ${\bm{x}}_{0}$ through a joint distribution $\pi_{\theta}({\bm{x}}_{t},{\bm{x}}_{0})=\pi_{\theta}({\bm{x}}_{0})\pi_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})$).
@@ 3.3 Practical Implementation
Adaptive Loss Weighting. Typical diffusion loss includes a time-dependent weighting $w(t)$ (Eq. (1)). Instead of manual tuning, we adopt an adaptive weighting scheme. The velocity predictor ${\bm{v}}_{\theta}$ can be equivalently transformed into ${\bm{x}}_{0}$ predictor, denoted as ${\bm{x}}_{\theta}$ (e.g., ${\bm{x}}_{\theta}={\bm{x}}_{t}-t{\bm{v}}_{\theta}$ under rectified flow schedule). We replace the weighting with a form of self-normalized ${\bm{x}}_{0}$ regression, motivated by the diffusion distillation method DMD (Yin et al., 2024):
@@
where sg is the stop-gradient operator. We find it typically leads to faster training (Figure 9).
CFG-Free Optimization. Classifier-Free Guidance (CFG) (Ho & Salimans, 2022) is a default technique to enhance generation quality at inference time, yet it complicates post-training and reduces efficiency. Conceptually, we interpret CFG as an offline form of reinforcement guidance (Eq. (3.1)), where conditional and unconditional models correspond to positive and negative signals. With this understanding, we discard CFG in our algorithm design, and the policy is initialized solely by the conditional model. Despite this seemingly poor initialization, we observe that performance surges and quickly surpasses the CFG baseline (Figure 1). This suggests that the functionality of CFG can be effectively learned or substituted through RL post-training, echoing recent studies that achieve strong performance without CFG through post-training (Chen et al., 2025b; a; Zheng et al., 2025).
@@ 5 Related Work
Likelihood-based methods: (1) Diffusion-DPO (Wallace et al., 2024; Yang et al., 2024; Liang et al., 2024; Yuan et al., 2024; Li et al., 2025a) adapts DPO to diffusion for paired human preference data but requires additional likelihood and loss approximations compared to AR; DDO (Zheng et al., 2025) uses high-quality dataset as positive signals and self-generated samples as negative signals to avoid the requirement of paired data, achieving state-of-the-art CFG-free FIDs in visual generation, while still relying on likelihood approximation for the diffusion case. (2) Policy gradient methods, starting from PPO style (Black et al., 2023; Fan et al., 2023), decompose trajectory likelihoods step by step without considering forward consistency. Recent GRPO extensions (Liu et al., 2025; Xue et al., 2025) prove effective and scalable for diffusion RL, but they couple the training loss with SDE samplers and face efficiency bottlenecks. MixGRPO (Li et al., 2025b) improves efficiency by mixing SDE and ODE, while issues of coupling and forward inconsistency remain.
@@ 6 Conclusion
We introduce Diffusion Negative-aware FineTuning (DiffusionNFT), a new paradigm for online reinforcement learning of diffusion models that directly operates on the forward process. By formulating policy improvement as a contrast between positive and negative generations, DiffusionNFT integrates reinforcement signals seamlessly into the standard diffusion objective, eliminating the reliance on likelihood estimation and SDE-based reverse process. Empirically, DiffusionNFT demonstrates strong and efficient reward optimization, achieving up to $25\times$ higher efficiency than FlowGRPO while producing a single model that outperforms CFG baselines across diverse in-domain and out-of-domain rewards. We believe this work represents a step toward unifying supervised and reinforcement learning in diffusion, and highlights the forward process as a promising foundation for scalable, efficient, and theoretically principled diffusion RL.
The Use of Large Language Models (LLMs)
We used large language models (LLMs) solely as a writing assistant for language polishing and improving clarity of presentation. The LLMs were not involved in research ideation, methodological design, experimental execution, or result analysis. All scientific contributions and substantive writing were carried out by the authors.
@@ B.1 Flow SDE
To leverage the diffusion SDE formulation in Song et al. (2020b), we need to match its forward SDE $\mathrm{d}{\bm{x}}_{t}=f(t){\bm{x}}_{t}\mathrm{d}t+g(t)\mathrm{d}{\bm{w}}_{t}$ with the forward transition kernel ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}\bm{\epsilon}$. As noted in the first two arXiv versions of the VDM paper (Kingma et al., 2021), $f(t),g(t)$ are related to $\alpha_{t},\sigma_{t}$ by $f(t)=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}$, $g^{2}(t)=\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma_{t}^{2}$. Setting $\alpha_{t}=1-t,\sigma_{t}=t$, we have
@@
where $\lambda_{t}\in[0,1]$. Equivalently, it amounts to introducing Langevin dynamics on top of the diffusion ODE, with $\lambda_{t}=0$ corresponding to ODE, and $\lambda_{t}=1$ corresponding to the maximum variance SDE in Song et al. (2020b). The score function ${\bm{s}}_{\theta}({\bm{x}}_{t},t)\approx\nabla_{{\bm{x}}_{t}}\log\pi_{t}({\bm{x}}_{t})$, noise predictor $\bm{\epsilon}_{\theta}({\bm{x}}_{t},t)$, data predictor ${\bm{x}}_{\theta}({\bm{x}}_{t},t)$ and velocity predictor ${\bm{v}}_{\theta}({\bm{x}}_{t},t)$ are interconvertible under general noise schedules (Zheng et al., 2023b):
@@
where $\rho_{t}=\eta_{t}s\sqrt{1-\frac{s^{2}(1-t)^{2}}{t^{2}(1-s)^{2}}}$, and $\eta_{t}\in[0,1]$ interpolates between ODE and maximum variance SDE. Compared to the Euler discretization, the DDIM-style discretization avoids singularities at boundaries and is expected to reduce sampling errors. However, we did not observe notable advantages by replacing the SDE sampler with stochastic DDIM. Concurrent work (Wang & Yu, 2025) improves the SDE sampler through the Coefficients-Preserving Sampling (CPS) principle.
@@ B.3 Intuition behind the FlowGRPO Objective
Therefore, FlowGRPO essentially aligns the velocity field with the advantage-weighted noise, while the choice of timesteps and sampler only influences the weighting $\frac{m(s,t)}{n(s,t)}$ across sampling steps. In the following, we show a further conclusion that FlowGRPO can be viewed as a gradient estimation of reward backpropagation.
Denote $r_{t}({\bm{x}}_{t})$ as the implicit gradient-free function that solves the PF-ODE from $t$ to 0 and fetches the reward on the cleaned sample. The rewards can be expressed as
$r^{(i)}=r_{s}\left(l(s,t){\bm{x}}_{t}-m(s,t){\bm{v}}_{\theta}({\bm{x}}_{t},t)+n(s,t)\bm{\epsilon}^{(i)}\right)$ (23)
According to Stein’s identity, we have
$\frac{1}{N}\sum_{i=1}^{N}r^{(i)}\bm{\epsilon}^{(i)}$ $\approx\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[r_{s}\left(l(s,t){\bm{x}}_{t}-m(s,t){\bm{v}}_{\theta}({\bm{x}}_{t},t)+n(s,t)\bm{\epsilon}\right)\bm{\epsilon}\right]$ (24)
$=n(s,t)\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\nabla r_{s}\left(l(s,t){\bm{x}}_{t}-m(s,t){\bm{v}}_{\theta}({\bm{x}}_{t},t)+n(s,t)\bm{\epsilon}\right)\right]$
Therefore,
$\nabla_{\theta}\left[\frac{1}{N}\sum_{i=1}^{N}(A^{(i)}\bm{\epsilon}^{(i)})^{\top}{\bm{v}}_{\theta}({\bm{x}}_{t},t)\right]$ (25)
$\approx$ $\frac{n(s,t)}{\sigma}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\nabla r_{s}\left(l(s,t){\bm{x}}_{t}-m(s,t){\bm{v}}_{\theta}({\bm{x}}_{t},t)+n(s,t)\bm{\epsilon}\right)\nabla_{\theta}{\bm{v}}_{\theta}({\bm{x}}_{t},t)\right]$
$=$ $-\frac{n(s,t)}{m(s,t)\sigma}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\nabla_{\theta}r_{s}\left(l(s,t){\bm{x}}_{t}-m(s,t){\bm{v}}_{\theta}({\bm{x}}_{t},t)+n(s,t)\bm{\epsilon}\right)\right]$
where $\sigma$ is the global std used in GRPO normalization. Therefore, the GRPO loss gradient is
$\nabla_{\theta}\mathcal{L}(\theta)\approx-\frac{1}{\sigma}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\nabla_{\theta}r_{s}\left(l(s,t){\bm{x}}_{t}-m(s,t){\bm{v}}_{\theta}({\bm{x}}_{t},t)+n(s,t)\bm{\epsilon}\right)\right]$ (26)
From the above gradient, GRPO optimizes the reverse transition $t\rightarrow s$ when the remaining trajectory $s\rightarrow 0$ is gradient-free. Compared to works like ReFL (Xu et al., 2023), which conduct direct gradient backpropagation and approximate $s\rightarrow 0$ with a single forward pass (${\bm{x}}_{0}$-prediction), GRPO introduces higher estimation variance but avoids backpropagation through the $s\rightarrow 0$ process, allowing larger $s$ and a longer sampling chain for $s\rightarrow 0$.